### v1

@@ 4.2.2 Projection Head
where $\alpha$ and $\beta$ denote the weights for the hidden layer and output layer, respectively. $\sigma(\cdot)$ is the ReLU activation function. Weights and bias terms are initialized. The initial values of the weight matrix follow the Kaiming initialization strategy, while those of the bias term follow a uniform distribution. The forward method encapsulates the logic of forward propagation, which involves multiplying the input data by a weight matrix plus a bias term.
@@ 4.2.3 Learnable Data Augmentation
The purpose of the augmentation is to preserve important features and filter out noisy data, which is the basis for the later selection of positive and negative pairs in contrastive learning. For the teacher network, we mask the high-dimensional mapping $\mathbf{z_{i}}$ after the Multi-Layer Perceptron (MLP); for the student network, we mask the original data $\mathbf{x_{i}}$ before the MLP. This approach is driven by performance considerations, and we present the detailed analysis in the following experimental section. The two augmentation strategies enhance the representation learning ability of the encoder and make them pay more attention to the valuable information in the time series rather than data noise. Specifically, we mask the pivotal timestamps of $\mathbf{x_{i}}$ to generate augmented views, which can be expressed as follows:
${g}_{TD}=\{\{x_{i,t}\odot\rho_{t}\mid x_{i,t}\in\mathcal{Z}\},\mathcal{E}\}.$ (3)
where $\rho_{t}\in\{0,1\}$ is drawn from a Bemoulli distribution parameterized by $\omega_{t}$, i.e., $\rho_{t}\in Bern(\omega_{t})$, which denotes whether to keep the observation $x_{i,t}$ of instance $\mathbf{x_{i}}$.
Afterwards, the augmented views are respectively fed into the teacher and student network to obtain the representations.
@@ 4.3.1 Momentum Encoder
The encoder is designed to extract rich information and generate representations from the augmented time series. Most existing networks can be applied in our framework, we choose dilated casual convolution and residual connections as the backbone of the encoder. The architecture of the encoder is illustrated in Fig.2. It stacks multiple layers to extract multiple temporal dependencies in two aspects (i) between different instances, and (ii) between different timestamps of the same instance. The representations after encoding flow in two ways, the representation of contrastive learning and the supervised task.
We first introduce the dilated causal convolution. Given a 1-D sequence input $\mathbf{z}\in\mathbb{R}^{n}$ and a function $f:\{0,\ldots,k-1\}\to\mathbb{R}$, we represent the causal convolution on element $s$ as:
@@
In order to achieve a larger receptive field for each input, multiple causal convolution layers are usually stacked. However, it may arise out of memory problem. In most existing works, dilated causal convolution utilized to provide the exponential expansion of the receptive field. Formally, the dilated convolution operation $F$ is represented as:
@@
where $d$ is the dilation factor, $k$ is the filter size, $d$ increases exponentially according the depth of the network. When $d=1$, the dilated convolution operator *$*d$ reduces to a regular convolution.
Fig.2 shows the backbone of the encoder, which follows the sequence of GELU $\to$ DilatedConv $\to$ GELU $\to$ DilatedConv [2]. Specifically, each DilatedConv unit comprises two $1-D$ convolutional layers with dilation parameter ($2^{p}$ is the dilation parameter of $p-$th block, $p\in\{1,2,\cdots,P\}$). The Gaussian Error Linear Unit (GELU)[13] is a powerful activation function defined as $x\Phi(x)$, where $\Phi(x)$ is the standard Gaussian cumulative distribution function. GELU is formulated as following,
@@
In addition, residual connections [12] between adjacent layers are employed to facilitate the model ability to acquire deeper semantic information. These components cooperate to joinlty extract the context representation for each timestamp.
@@ 4.4 Momentum Contrastive Learning Task
where momentum coefficient $m\in[0,1)$. As shown in Fig.1, back-propagation is employed to update $\theta_{s}$, while Eq.(11) is used to update $\theta_{t}$.
@@ 4.5 Adaptive Supervised Task
Finally, we employ the prevalent cross-entropy loss[8] as the supervised loss function to control the optimization of the module. The objective of the cross-entropy loss is to compute the similarity $s(h_{t},h_{s})$ easily between two probability matrices, the two probabilities $p_{i}^{s}$ and $p_{i}^{t}$ after softmax correspond to the two probability distributions in the cross-entropy loss, which is equivalent to doing MLE (Maximum Likelihood Estimation) on $p_{i}^{s}$ and $p_{i}^{t}$, i.e., making them more similar, consistent with our conceptualization of evaluating not only the distances between positive and negative samples within the sample space, but also the comparability of their respective pairs. It is formulated as:
@@
4.6 Joint Optimization and Prediction
where $\lambda$ is a hyperparameter representation the the relative weight of the supervised task. The experimental results have shown that this joint optimization is capable of improving the model’s representation learning capabilities and boosting its performance in downstream tasks.
where $\hat{\mathbf{y_{i}}}$ is the predicted value, $\mathbf{y_{i}}$ is the ground truth. Then, we train the model using Mean Absolute Error (MAE) as the regression loss. Ridge regression is a regularization technique commonly employed in regression tasks, including time series forecasting. It introduces a penalty term to the loss function, which helps to mitigate overfitting and stabilize the model’s predictions. In our model, we use cross-validation to identify the optimal ridge regression model for the forecasting task, which is then used in subsequent forecasting stage.
@@ 4.6 Joint Optimization and Prediction
Algorithm 1 DE-TSMCL Pre-train PyTorch-like style pseudocode.
@@ 5 Experimental Results and Analysis
  • RQ4. How dose the learned data augmentation influence the the model performance?
@@ 5.1.1 Dataset Description
We conduct experiments on four public real-world datasets, including ETTh1, ETTh2, ETTm1 and Electricity. The summarized statistics of the datasets are presented in Table 2. Following existing works[37, 31], we partition the data into training, validation testing set with a ratio of 3:1:1 in the experiments.
  • ETT(Electricity Transformer Temperature)[42]: The ETT serves as a key indicator in the long-term deployment of electric power. This dataset encompasses data spanning two years from two distinct counties in China. In order to delve into the granularity of the Long Sequence Time-Series Forecasting (LSTF) issue, various subsets were formulated: ETTh1, ETTh2 for the 1-hour level, and ETTm1 for the 15-minute level. Each data point comprises the target value ‘oil temperature’ along with six power load features. The duration for training/validation/testing is segmented into 12/4/4 months respectively.
  • Electricity[32] : The electricity dataset is thatelectric power consumption measurements were taken in one household at a one-minute sampling rate over a nearly 4-year period. Various electrical quantities and some sub-metering values are available. This dataset contains 2,075,259 measurements collected in a house located in Sceaux, France (7 km from Paris) between December 2006 and November 2010 (spanning 47 months).
@@ 5.1.2 Compared Methods
  • N-BEATS [22]: Proposes a block-based network structure to learn intricate timing dependencies. The prediction results from various modules are integrated through combination operations to accomplish multi-view time series modeling.
  • LogTrans [16]: Utilizing log-transformed series data for improved time series predictions. Address the issue of non-static trends present in the original data.
  • Informer [42]: Presenting ProbSparse self-attention mechanism enhanced transformer architecture for time series forecasting. It mitigates the complexity of self-attention by employing random sampling and probability distribution, This method can handle extremely lengthy time series. Furthermore, its integration with the local self-attention block allows for simultaneous modeling of global dependencies and local trends.
@@ 5.1.3 Evaluation Metrics
$\mathrm{MSE}=\frac{1}{HC}\sum_{i=1}^{H}\sum_{j=1}^{C}\left(x_{t+i}^{(j)}-\hat{x}_{t+i}^{(j)}\right)^{2},$ (17)
$\mathrm{MAE}=\frac{1}{HC}\sum_{i=1}^{H}\sum_{j=1}^{C}|x_{t+i}^{(j)}-\hat{x}_{t+i}^{(j)}|.$ (18)
@@ 5.1.4 Training Details &\&& Hyperparameters
The model has been implemented based on the PyTorch framework. All experiments have been carried out on a Intel(R) Core i5-8400 @@@@ 2.80GHz hardware platform equipped with NVIDIA GeForce GTX 1060(6G) GPU. Specially, the default batch size is 4, and the learning rate is $1e^{-3}$. For ETTh1 and ETTh2 datasets, the number of pre-training iterations is 200, while for ETTm1 and electricity datasets, it is increased to 600. The dimension of representations is 320. The channels of MLP projection head hidden layers is 64. Following existing works such as Ts2vec[37] and Informer[42], the lengths of look-back window is $l$in {24, 48, 96, 288, 672} for the ETTm1 dataset and $l$in {24, 48, 168, 336, 720} for the remaining datasets.
@@ 5.2 Main Results(RQ1)
TABLE 3 and TABLE 4 report the comparison results for univariate time series and multivariate time series forecasting, respectively. We will analyze the experimental results in the rest of the section.
@@ 5.2.1 Univariate Time Series Forecasting
The results for univariate time series forecasting of ten different methods for four datasets are tabulated in Table3. On all the different datasets, DE-TSMCL consistently obtains the best performance under two evaluation metrics. In particular, when compared with the most up-to-date contrastive learning based method TS2Vec, DE-TSMCL achieves obvious performance improvement. For example, DE-TSMCL achieves 24.2% improvement in MSE and 14.7% in MAE on the ETTm1 dataset, which contains a considerable volume of data. On the ETTh1 dataset, DE-TSMCL also gets similar gains(14.2% improvement in MSE and 10.6% in MAE). Such significant performance gains are primarily attributed to momentum update and distillation enhanced learning framework. A larger dataset can benefit the training process of both the teacher and student networks in a knowledge distillation setup. When the dataset constains a large volume of data, it allows the teacher network to undergo more updates, refining its knowledge and providing a more accurate target distribution for the student. Thus, the joint training of the teacher and student networks can lead to improved representation and prediction results by leveraging the teacher’s expertise and the diversity of the data.
@@ 5.2.2 Multivariate Time Series Forecasting
The results for multivariate time series forecasting of nine methods for four datasets are reported in Table 4. It is easy to observe that the proposed DE-TSMCL model consistently outperforms other state-of-the-art approaches on all datasets. These results indicate that the pre-trained encoder effectively captures the features of time series to improve prediction performance. Similar to univariate time series forecasting, DE-TSMCL achieves remarkable improvements on ETTm1 and electricity datasets. Specifically, we achieve a maximum improvement of 12.2% on the MSE criterion and 27.3% on the MAE criterion. In addition, we note that the improvement in multivariate forecasting is superior to that in univariate time series forecasting, which is consistent with our findings in the previous section indicating that momentum updating benefits more from more training data.
Table 5: Ablation results with look-back window $l\in\{24,68,168,336,720\}$ for ETTh1. Results are averaged from all prediction lengths.
@@
Table 6: Ablation results with look-back window $l\in\{24,68,168,336,720\}$ for ETTh2. Results are averaged from all prediction lengths.
@@ 5.3 Ablation Study(RQ2)
TABLE 5 and TABLE 6 report the experimental results of all the variants for ETTh1 and ETTh2 datasets, respectively. We start from the basic forecasting framework and introduce each component or their combination in turn. As shown in the table, adopting distillation enhanced framework(DE+M) has provided 2.9% (MSE) and +2.1% (MAE) improvement for univariate time series forecasting, as well as 0.4% (MSE) and 0.2% (MAE) improvement for multivarite time series forecasting on ETTH1. For ETTH2 dataset, the improvement is up to 7.3% (MSE) and 2.4% (MAE) for multivarite time series forecasting. Moreover, when combing with other components, the improvement can increase to 7.1% (DE+M+C), 7.5% (DE+M+S) and 7.7% (DE-TSMCL). Theses results demonstrate that our proposed distillation enhanced framework with with momentum update is able to obviously improve the representation capability, moreover help achieve better forecasting performance than the basic time series prediction method.
In order to visualize the impact of each component on all look-back windows, we plot Fig. 3 and Fig. 4 to visually present the forecasting results of four basic variants for ETTm1 and electricity datasets. We can observe that the performance improves by incorporating distillation framework with momentum update, supervised task and center layer, which demonstrates that the designed framework is effective to enhance the representation ability of the model to improve the prediction performance. First, it is obvious that the performance of Basic/Basic+C/Basic+S is worse than DE-TSMCL, indicating the advantage of utilizing the distillation framework with momentum update for learning representations of time series data. Moreover, the advantage exhibits variations across different datasets. The main reason is that ETTm1 contains greater volume of data, which achieves more benefit from the training process of both the teacher and student networks in a knowledge distillation setup. Second, the performance of DE+M is also inferior to DE-TSMCL, which invalidates the function of supervised task. In addition, it is observed that the inclusion of only center layer (Basic+C) or supervised task(Basic+S) also slightly improves the performance of our model. Therefore, the superior performance of DE-TSMCL demonstrates the effectiveness of our design in introducing distillation framework with momentum update and supervised task for time series forecasting.
Refer to caption (a) MSE on ETTm1
Refer to caption (b) MAE on ETTm1
Refer to caption (c) MSE on Electricity
Refer to caption (d) MAE on Electricity
@@
Refer to caption (a) MSE on ETTm1
Refer to caption (b) MAE on ETTm1
Refer to caption (c) MSE on Electricity
Refer to caption (d) MAE on Electricity

### v2

@@ 4.2.2 Projection Head
where $\alpha$ and $\beta$ denote the weights for the hidden layer and output layer, respectively. $\sigma(\cdot)$ is the ReLU activation function. Weights and bias terms are initialized. The initial values of the weight matrix follow the Kaiming initialization strategy, while those of the bias terms follow a uniform distribution. The forward method encapsulates the logic of forward propagation, which involves multiplying the input data by a weight matrix plus a bias term.
@@ 4.2.3 Learnable Data Augmentation
The purpose of the augmentation is to preserve important features and filter out noisy data, which is the basis for the later selection of positive and negative pairs in contrastive learning. For the teacher network, we mask the high-dimensional mapping $\mathbf{z_{i}}$ after the MLP. For the student network, we mask the original data $\mathbf{x_{i}}$ before the MLP. This approach is driven by performance considerations, and we present the detailed analysis in the following experimental section. The two augmentation strategies enhance the representation learning ability of the encoder and make them pay more attention to the valuable information in the time series rather than data noise. Specifically, we mask the pivotal timestamps of $\mathbf{x_{i}}$ to generate augmented views, which can be expressed as follows:
${g}_{TD}=\{\{x_{i,t}\odot\rho_{t}\mid x_{i,t}\in\mathcal{X}\},\mathcal{E}\}.$ (3)
where $\rho_{t}\in\{0,1\}$ is drawn from a Bemoulli distribution parameterized by $\omega_{t}$, i.e., $\rho_{t}\in Bern(\omega_{t})$, which denotes whether to keep the observation $x_{i,t}$ of instance $\mathbf{x_{i}}$. Afterwards, the augmented views are respectively fed into the teacher and student networks to obtain the representations.
@@ 4.3.1 Momentum Encoder
The encoder is designed to extract rich information and generate representations from the augmented time series. Most existing networks can be applied in our framework, we choose dilated casual convolution and residual connections as the backbone of the encoder. The architecture of the encoder is shown in Fig.2. It stacks multiple layers to extract multiple temporal dependencies in two aspects (i) between different instances, and (ii) between different timestamps of the same instance. The representations after encoding flow in two ways, the contrastive learning representation and the supervised task representation.
We first introduce the dilated causal convolution. Given a 1-D sequence input $\mathbf{z}\in{R}^{n}$ and a function $f:\{0,\ldots,k-1\}\to{R}$, we represent the causal convolution on element $s$ as:
@@
To achieve a larger receptive field for each input, multiple causal convolution layers are usually stacked, so in this paper we use $Q$ as the maximum stacked convolution layers. In most existing works, dilated causal convolution is used to provide the exponential expansion of the receptive field. Formally, the dilated convolution operation $F$ is represented as
@@
where $d$ is the dilation factor, $k$ is the filter size, $d$ increases exponentially according the depth of the network. When $d=1$, the dilated convolution operator $*d$ reduces to a regular convolution.
Fig.2 shows the backbone of the encoder, which follows the sequence of GELU $\to$ DilatedConv $\to$ GELU $\to$ DilatedConv. Specifically, each DilatedConv unit comprises two $1-D$ convolutional layers with dilation parameter ($2^{p}$ is the dilation parameter of $p-$th block, $p\in\{1,2,\cdots,Q\}$). The Gaussian Error Linear Unit (GELU)[14] is a powerful activation function defined as $x\Phi(x)$, where $\Phi(x)$ is the standard Gaussian cumulative distribution function. GELU is formulated as following,
@@
In addition, residual connections [13] between adjacent layers are employed to facilitate the model ability to acquire deeper semantic information. These components cooperate to jointly extract the context representation for each timestamp.
In addition to the TCN[1] used in our encoder, we can incorporate other convolutional frameworks into our model, such as sparse convolutions [2, 28], which have shown promise in the field of 3D reconstruction. Notably, the study presented in [2] introduces the concept of sparsity-invariant convolution for handling irregularly sampled time series. This research offers valuable insights that can significantly contribute to solving the challenge of forecasting irregularly sampled time series using our proposed model.
@@ 4.4 Momentum Contrastive Learning Task
where momentum coefficient $m\in[0,1)$. As shown in Fig.1, back-propagation is employed to update $\theta_{s}$, while Eq.(11) is used to update $\theta_{t}$. Inspired by the work of DINO[3], we utilize the stop gradient operator (s-g) on the teacher network to restrict the propagation of gradients exclusively through the student network. This ensures that during the back-propagation process, the gradients are not updated for the teacher network. Instead, the teacher parameters are updated using the exponential moving average (Ema) of the student parameters. Conversely, for the student network, the gradients are updated normally, allowing for the learning and optimization of its parameters. To maintain consistency and visual clarity in representing the flow of gradients within the teacher-student network, we also include a gradient return marker of the same color for the student network. This decision aligns with our pseudocode and promotes a uniform representation of operators throughout the network architecture.
@@ 4.5 Adaptive Supervised Task
Finally, we employ the prevalent cross-entropy loss[9] as the supervised loss function to control the optimization of the module. The objective of the cross-entropy loss is to compute the similarity $s(h_{t},h_{s})$ between two probability matrices, the two probabilities $p_{i}^{s}$ and $p_{i}^{t}$ after softmax correspond to the two probability distributions in the cross-entropy loss, which is equivalent to implementing MLE (Maximum Likelihood Estimation) on $p_{i}^{s}$ and $p_{i}^{t}$, i.e., making them more similar, consistent with our conceptualization of evaluating not only the distances between positive and negative samples within the sample space, but also the comparability of their respective pairs. It is formulated as:
@@
Algorithm 1 DE-TSMCL Pre-train PyTorch-like style pseudocode.
@@
4.6 Joint Optimization and Prediction
where $\lambda$ is a hyperparameter representing the relative weight of the supervised task. The experimental results have shown that this joint optimization is capable of improving the model’s representation learning capabilities and boosting its performance in downstream tasks.
where $\hat{\mathbf{y_{i}}}$ is the predicted value, $\mathbf{y_{i}}$ is the ground truth. Then, we train the model using Mean Absolute Error (MAE) as the regression loss. Ridge regression is a regularization technique commonly employed in regression tasks, including time series forecasting. It introduces a penalty term to the loss function, which helps mitigate overfitting and stabilize the model’s predictions. In our model, we use cross-validation to identify the optimal ridge regression model for the forecasting task, which is then used in subsequent forecasting stage.
An overview of the training process for DE-TSMCL is provided in Algorithm 1. DE-TSMCL is an efficient algorithm when compared to traditional contrastive learning models. The complexity of DE-TSMCL is $O(TCK)+O(T)+O(KL)+O(KL^{2})$, where $O(TCK)$ denotes the complexity of the projection layer, $O(T)$ is the complexity of the data augmentation module, the complexity of the momentum encoder and supervised loss is $O(KL)$, and $O(KL^{2})$ represents the complexity of the InfoNCE loss. In DE-TSMCL, the sub-series length $T$ scales linearly with both the projection layer and augmentations. The joint optimization module scales quadratically with the number of overlapping sub-series $L$ and linearly with the dimension $K$.
@@ 5 Experimental Results and Analysis
  • RQ4. How does the learned data augmentation influence the the model performance?
@@ 5.1.1 Dataset Description
We conduct experiments on five public real-world datasets, including ETTh1, ETTh2, ETTm1, ETTm2 and Electricity. The summarized statistics of the datasets are presented in Table 2. Following existing works[42, 33], we partition the data into training, validation testing set with a ratio of 3:1:1 in the experiments.
  • ETT(Electricity Transformer Temperature)[47]: The ETT serves as a key indicator in the long-term deployment of electric power. This dataset encompasses data spanning two years from two distinct counties in China. In order to delve into the granularity of the long sequence time-series forecasting issue, various subsets are formulated: ETTh1, ETTh2 for the 1-hour level, and ETTm1,ETTm2 for the 15-minute level. Each data point comprises the target value "oil temperature" along with six power load features.
  • Electricity[34]: The electricity dataset is the electric power consumption measurements are taken in one household at a one-minute sampling rate over a nearly 4-year period. Various electrical quantities and some sub-metering values are available. This dataset contains 2,075,259 measurements collected in a house located in Sceaux, France (7 km from Paris) between December 2006 and November 2010 (spanning 47 months).
@@ 5.1.2 Compared Methods
  • N-BEATS [23]: It is a block-based network structure to learn intricate timing dependencies. The prediction results from various modules are integrated through combination operations to accomplish multi-view time series modeling.
  • LogTrans [18]: It utilizes log-transformed series data for improved time series predictions. Address the issue of non-static trends present in the original data.
  • Reformer [16]: It introduces local-sensitive hashing (LSH) to approximate attention by allocating similar queries. Not only are subsequent models improved by reduced complexity, but they also further develop intricate building blocks for time series forecasting.
  • Autoformer [34]: It utilizes autocorrelation to establish patch-level connections, but this is a manually designed approach that does not encapsulate all semantic information within a patch.
  • Informer [47]: It presents an enhanced transformer architecture for time series forecasting. It mitigates the complexity of self-attention by employing random sampling and probability distribution.
@@ 5.1.3 Evaluation Metrics
$\mathrm{MSE}=\frac{1}{PC}\sum_{i=1}^{P}\sum_{j=1}^{C}\left(x_{t+i}^{(j)}-\hat{x}_{t+i}^{(j)}\right)^{2},$ (17)
$\mathrm{MAE}=\frac{1}{PC}\sum_{i=1}^{P}\sum_{j=1}^{C}|x_{t+i}^{(j)}-\hat{x}_{t+i}^{(j)}|.$ (18)
@@ 5.1.4 Training Details &\&& Hyperparameters
The model has been implemented based on the PyTorch framework. All experiments have been carried out on a Intel(R) Core i5-8400 @@@@ 2.80GHz hardware platform equipped with NVIDIA GeForce GTX 1060(6G) GPU. Specially, the default batch size is 4, and the learning rate is $1e^{-3}$. For ETTh1 and ETTh2 datasets, the number of pre-training iterations is 200, while for ETTm1, ETTm2 and electricity datasets, it is increased to 600. The dimension of representations is 320. The channels of MLP projection head hidden layers is 64. Following existing works such as TS2vec[42] and Informer[47], the prediction length is $P$in {24, 48, 96, 288, 672} for the ETTm1 dataset and $P$in {24, 48, 168, 336, 720} for the remaining datasets.
@@ 5.2 Main Results(RQ1)
Table 3 and Table 4 report the comparison results of different approaches for univariate time series and multivariate time series forecasting, respectively. We will analyze the experimental results in the rest of the section.
Table 5: Comparison results of different approach for univariate and multivariate forecasting on ETTm2 dataset.
P DE-TSMCL Informer LogTrans Reformer
MSE MAE MSE MAE MSE MAE MSE MAE
Univariate 96 0.086 0.199 0.088 0.225 0.075 0.208 0.131 0.288
192 0.118 0.251 0.132 0.283 0.129 0.275 0.186 0.354
336 0.152 0.301 0.180 0.336 0.154 0.302 0.220 0.381
720 0.200 0.319 0.300 0.435 0.160 0.321 0.267 0.430
Multivariate 96 0.304 0.356 0.355 0.462 0.768 0.642 0.658 0.619
192 0.315 0.369 0.595 0.586 0.989 0.757 1.078 0.827
336 0.337 0.381 1.270 0.871 1.334 0.872 1.549 0.972
720 0.398 0.415 3.001 1.267 3.048 1.328 2.631 1.242
@@ 5.2.1 Univariate Time Series Forecasting
The results for univariate time series forecasting of ten different methods for four datasets are tabulated in Table3. On all the different datasets, DE-TSMCL consistently obtains the best performance under two evaluation metrics. In particular, when compared with the up-to-date contrastive learning based method TS2Vec, DE-TSMCL achieves obvious performance improvement. For example, DE-TSMCL achieves 24.2% improvement in MSE and 14.7% in MAE on the ETTm1 dataset, which contains a considerable volume of data. On the ETTh1 dataset, DE-TSMCL also gets similar gains(14.2% improvement in MSE and 10.6% in MAE). Such significant performance gains are primarily attributed to momentum update and distillation enhanced learning framework. A larger dataset can benefit the training process of both the teacher and student networks in a knowledge distillation setup. When the dataset constains a large volume of data, it allows the teacher network to undergo more updates, refining its knowledge and providing a more accurate target distribution for the student. Thus, the joint training of the teacher and student networks can lead to improved representation and prediction results by leveraging the teacher’s expertise and the diversity of the data.
@@ 5.2.2 Multivariate Time Series Forecasting
The results for multivariate time series forecasting of nine methods for four datasets are reported in Table 4. It is easy to observe that the proposed DE-TSMCL model consistently outperforms other state-of-the-art approaches on all datasets. These results indicate that the pre-trained encoder effectively captures the features of time series to improve prediction performance. Similar to univariate time series forecasting, DE-TSMCL achieves remarkable improvements on ETTm1 and electricity datasets. Specifically, we achieve a maximum improvement of 12.2% on the MSE criterion and 27.3% on the MAE criterion.
To provide additional experimental results in the energy domain, we conduct experiments for both univariate and multivariate forecasting using the ETTm2 dataset. Table 5 presents the comparison results of different approaches. We observe that our model consistently demonstrates improved performance as the prediction length increases. Specifically, the accuracy increased from 3.3‰to 6.2‰when the prediction length extends from 336 to 720 time steps for univariate forecasting. This indicates that our model becomes more accurate with longer prediction horizons. Furthermore, we note that the improvement in multivariate forecasting is greater compared to univariate time series forecasting. This observation aligns with the findings from the previous section, which indicates that the momentum updating technique benefits more from having more training data.
Table 6: Ablation results with prediction lengths $P\in\{24,68,168,336,720\}$ for ETTh1. Results are averaged from all prediction lengths.
@@
Table 7: Ablation results with prediction lengths $P\in\{24,68,168,336,720\}$ for ETTh2. Results are averaged from all prediction lengths.
@@ 5.3 Ablation Study(RQ2)
Table 6 and Table 7 report the experimental results of all variants for ETTh1 and ETTh2 datasets, respectively. We start from the basic forecasting framework and introduce each component or their combination in turn. As shown in the table, adopting distillation enhanced framework(DE+M) has provided 2.9% (MSE) and +2.1% (MAE) improvement for univariate time series forecasting, as well as 0.4% (MSE) and 0.2% (MAE) improvement for multivarite time series forecasting on ETTh1. For ETTh2 dataset, the improvement is up to 7.3% (MSE) and 2.4% (MAE) for multivarite time series forecasting. Moreover, when combing with other components, the improvement can increase to 7.1% (DE+M+C), 7.5% (DE+M+S) and 7.7% (DE-TSMCL). Theses results demonstrate that our proposed distillation enhanced framework with momentum update is able to obviously improve the representation capability, moreover help achieve better forecasting performance than the basic time series prediction method.
In order to visualize the impact of each component on all prediction , we plot Fig. 3 and Fig. 4 to visually present the forecasting results of four basic variants for ETTm1 and electricity datasets. We can observe that the performance improves by incorporating distillation framework with momentum update, supervised task and center layer, which demonstrates that the designed framework is effective to enhance the representation ability of the model to improve the prediction performance. First, it is obvious that the performance of Basic/Basic+C/Basic+S is worse than DE-TSMCL, indicating the advantage of utilizing the distillation framework with momentum update for learning representations of time series data. Moreover, the advantage exhibits variations across different datasets. The main reason is that ETTm1 contains greater volume of data, which achieves more benefit from the training process of both the teacher and student networks in a knowledge distillation setup. Second, the performance of DE+M is also inferior to DE-TSMCL, which invalidates the function of supervised task. In addition, it is observed that the inclusion of only center layer (Basic+C) or supervised task(Basic+S) also slightly improves the performance of our model. Therefore, the superior performance of DE-TSMCL demonstrates the effectiveness of our design in introducing distillation framework with momentum update and supervised task for time series forecasting.
(a) MSE on ETTm1
(b) MAE on ETTm1
(c) MSE on Electricity
(d) MAE on Electricity
@@
(a) MSE on ETTm1
(b) MAE on ETTm1
(c) MSE on Electricity
(d) MAE on Electricity